OpenStudy (amypurplediaz28):

I WILL FAN AND MEDAL!!!!!!!!!!!!!!!! Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%. Roger has 40 cars in his collection. Each year he collects 1 additional car. Part C: After approximately how many years is the number of cars that Cory and Roger have the same? Justify your answer mathematically. (4 points)

undeadknight26 (undeadknight26):

Well 20% of 15 is what.

OpenStudy (sshayer):

let x be the number of years then \[15(1+\frac{ 20 }{ 100 })^x=40 +1x\] \[\left( \frac{ 6 }{ 5 } \right)^x=\frac{ 40 }{ 15 }+\frac{ x }{ 15 }\] find x

OpenStudy (sshayer):

by hit and trial method after 6 years number of cars is almost the same.